what is the heat when 100.0g of mg cool from 725k ti 552k ?
Answer #1

Heat energy is given as,

Q = mc∆T

m -- mass of substance

C -- specific heat capacity

∆T = change in temperature

Specific heat of Mg = aaa 1.020 j/g °C

To convert temperature from Kelvin to degree Celsius , subtract 273 from it.

725k = 725-273. = 452 °C

552k = 552-273. = 279 °C

m = 100 g

Heat absorbed is ,

Q = 100 × 1.020 × (279 - 452)

Q = 100 × 1.020 × (-173)

Q = -17646 joules.

Negative sign implies that heat is absorbed.

Answer = 17646 joules